# GestureCoach: Rehearsing for Engaging Talks with LLM-Driven Gesture Recommendations

### v1

@@ Abstract.
Figure 1. GestureCoach guides speakers to perform gestures while rehearsing their talk. A gesture recommendation model predicts text segments in the presenter notes that should be emphasized with gestures and retrieves relevant semantic gestures for each segment. During rehearsal, the system highlights the segments and proactively cues a video clip of the gesture by tracking users’ speech, allowing them to integrate the gesture smoothly into their talk. Hovering over a segment allows users to preview and modify the associated gesture with alternate suggestions from the model.
[…]
The left side of the figure shows a speaker rehearsing for a presentation talk with the GestureCoach system. The speaker is saying ”one key reason” and showing the gesture symbolizing one by following a gesture cue provided by the system where a humanoid avatar demonstrates the same gesture. The right side shows a zoomed in view of the rehearsal interface. The interface has the appeal of the presenter view mode typically found in presentation software. It has 4 main windows. 1) the current slide 2) the presenter notes for the slide with text segments where the speaker should perform gestures highlighted in yellow. This window also has a rehearse button that speakers click to start rehearsal. 3) a window where a humanoid avatar displays the gesture that should be performed for each highlighted region. 4) a popup window which appears when speakers hover over a highlighted region. This window shows video clips of the top-3 gesture recommendations provided by the recommendation model as a carousel design. It has an edit and delete button that allows users to browse and select the gesture they prefer for a region or remove the region if they do not prefer to gesture there.
@@ 1. Introduction
As a step towards tackling these challenges, we propose GestureCoach, the first system capable of providing targeted guidance to perform gestures during rehearsal. GestureCoach combines a gesture recommendation system with a rehearsal interface that delivers real-time gesture cues, helping speakers naturally integrate gestures while speaking. The gesture recommendation system utilizes recent advances in AI chaining, by combining two specialized Large Language Model (LLM) modules, each targeting a specific aspect of gesture-based emphasis. First, the emphasis proposal module takes presenter notes as input and determines appropriate gesture regions. For this, we leverage the foundational knowledge of LLMs and further fine-tune it with annotated data from 10 TED talk videos of expert speakers. Second, the gesture identification module selects the specific semantic gesture to emphasize each predicted gesture region, by retrieving relevant examples from a curated database of gestures performed by experts using a Retrieval-Augmented Generation (RAG) framework  (Lewis et al., 2020).
@@ 2. Related Work
Figure 2. Overview of the GestureCoach system. Users’ presenter notes are processed by a gesture recommendation model in the backend. A frontend rehearsal interface displays the recommended gesture regions and gestures. During rehearsal, users’ speech is tracked and gesture cues are proactively delivered to provide real-time gesture guidance.
[…]
The figure shows a diagrammatic overview of the various components in the GestureCoach system and how they connect together. The presenter notes provided by the user is analysed by the Gesture recommendation model and the predicted gesture regions and gesture cues are passed to the rehearsal interface. During rehearsal, the interface tracks users’ speech to monitor their progress in the presenter notes and real-time gesture cues are fed back to user so that they can integrate them in their talk.
@@ 2.1. Co-speech Gestures for Communication
Co-speech gestures are hand and arm movements that naturally accompany speech and play a vital role in communication (Kendon, 1972; McNeill, 1992; Shattuck-Hufnagel and Prieto, 2019; Maricchiolo et al., 2020). These gestures are broadly classified into four types: beat, iconic, metaphoric, and deictic gestures (McNeill, 1992). Beat gestures are rhythmic movements that align with speech prosody to emphasize its structure. In contrast, iconic, metaphoric, and deictic gestures are semantic gestures that are directly related to the linguistic content being spoken. Iconic gestures physically resemble the objects or actions they represent, such as mimicking an object’s shape or reenacting a movement. Metaphoric gestures convey abstract concepts through spatial metaphors, like cupped hands symbolizing ”holding” an idea. Deictic gestures involve pointing to indicate referents, whether physically, such as gesturing toward an object, or abstractly, such as indicating an imaginative space.
Semantic gestures, which are the focus of this study, are context-driven and play a key role in enhancing content delivery during communication  (Maricchiolo et al., 2020; Kendon, 1994; Hostetter, 2011). Empirical evidence shows that gestures accompanying speech significantly benefit listener comprehension, especially when they depict motor actions or convey non-redundant information (Hostetter, 2011). These gestures also help speakers express their intentions through complementary movements, making the message clearer and more engaging for listeners (McNeill, 1992; Kendon, 2004; Hinnell, 2019). For example, moving one’s hands from side to side can indicate contrast when saying “on the other hand” (Hinnell, 2019). Moreover, encouraging speakers to use hand gestures has been found to increase narrative length and influence prosodic features, such as fundamental frequency and intensity, thereby improving intonation (Cravotta et al., 2019). These findings show that semantic gestures play an important role in making communication and presentations more effective.
@@ 2.3. Modeling Semantic Gestures
In this work, we aim to disentangle the tasks of predicting when to gesture and what semantic gesture to perform (Neff, 2016). Linguistic research suggests that semantic gestures often co-occur with specific spoken discourse markers, such as discourse connectives (e.g., “however,” “for example”) that indicate topic shifts, and stance markers (e.g., “absolutely” “probably”) that reflect the speaker’s level of certainty (Mcneill et al., 2014; Laparle, 2021; Calbris, 2011; Suresh et al., 2025). For instance, raised index fingers tend to co-occur with contrastive discourse markers like ”but,” ”however,” that signal exceptions (Inbar, 2022). These insights suggest that there are detectable patterns that can help determine when to gesture directly from text, which we address by fine-tuning an LLM to predict these gesture regions. To determine what semantic gesture to perform, i.e., the motion associated with a given context—we adopt a RAG-based approach, leveraging data from expert speakers to dynamically retrieve contextually appropriate gestures without being confined to a limited, predefined gesture set.
@@ 4. GestureCoach: Recommendation Model
Figure 3. Gesture recommendation model architecture. It consists of two LLM-based modules: 1) The emphasis proposal module is fine-tuned on data from expert speakers to predict gesture regions in the presenter notes. The predicted regions are filtered and passed to 2) the gesture identification module which selects the most suitable semantic gesture from the gesture database using Retrieval-Augmented Generation.
[…]
A diagrammatic figure detailing the architecture of the gesture recommendation model. The left side shows the flow for the emphasis proposal module. A sample presenter notes is shown and is provided as input to this module. The module predicts gesture regions in the notes as output. These regions are filtered and passed to the gesture identification module shown on the right side. For each predicted gesture region, a vector embedding is created using an embedding model. The embedding is used to query a gesture database that contains data from expert gesturing patterns. The top-k most similar gesture candidates from the database is retrieved and passed to an LLM, along with the predicted gesture region and the notes to select the most suitable gesture for the region from the candidates.
@@ 4.1. Gesture Database
To create a gesture database, we first collected TED talk videos of expert speakers who performed semantic gestures. The database consists of annotated gesture regions in the talk transcripts where the speaker performed semantic gestures. Each gesture region is also associated with a video clip of a humanoid avatar performing the semantic gesture. Details regarding the recording of these clips can be found in Section  6.
@@ 4.2.1. Fine-tuning with Data Augmentation
We include a filtering module that ensures predictions from the emphasis proposal module align with the original presenter notes before passing them to the gesture identification module. This step is important because the generative nature of LLMs can lead to predictions where the gesture regions do not match the original text verbatim. Such differences can range from minor ones (e.g., a missing apostrophe) to complete paraphrasing. The filtering process first attempts a verbatim match between the predicted gesture region and the presenter notes. If no exact match is found, we apply a sliding window approach to search for the most similar gesture region in the notes. Specifically, we compute the cosine similarity between the embedding of the predicted gesture region and the embeddings of each text window in the presenter notes. We set a similarity threshold of 75%, above which a match is accepted. If no sufficiently similar match is found, the prediction is discarded.
@@ 5. GestureCoach: Rehearsal Interface
Figure 4. Early Wizard-of-Oz prototype that simulates a real-time gesture cueing interface for rehearsing talks.
[…]
This figure visualises the setup for the early wizard-of-oz prototype based pilot study. A wizard is shown performing a gesture in front of a webcam. The speaker sees the demonstrated gesture in near real-time on a early prototype of the rehearsal interface. The speaker mimics the gesture as when the wizard shows a demonstration. A zoomed in view of the early prototype is also shown consisting of the highlighted presenter notes and view of the speaker from the webcam.
@@ 5.2. Pilot Study
We recruited 5 participants (3 males, 2 females, aged 25 to 45 years; among them four novice speakers and 1 expert speaker) who volunteered for the study. After an introduction and walkthrough of the prototype, participants were asked to rehearse for a talk, emphasizing the highlighted regions in the presenter notes by mimicking gesture cues displayed in the gesture window.
@@ 5.3. Design Insights & Final Interface
  • Gesture cue visualization Our pilots indicate that users were generally comfortable following gesture demonstrations presented in a third-person, upper-body view. This perspective allowed them to clearly observe arm and hand movements while understanding their spatial relation to the upper body, aiding effective gesturing. However, two participants reported that guidance from a real human felt distracting, inducing a sense of social evaluation, particularly due to eye contact and facial expressions: ”Its a bit embarrassing …I feel like someone’s looking at me… it make me more self-aware and affects my practice” (P2). Hence, in our final interface we use a humanoid avatar (upper-body only) captured from below the face, focusing on arm and hand motions to emphasize the gesture
@@
  • Modifying predicted gestures Users occasionally found that some of the gesture regions or the suggested gesture did not align with their expectations due to individual communication styles or cultural differences in gesturing norms. Certain gestures may carry different meanings across cultures, making them less suitable or even inappropriate for some users: ”I don’t point like this [pointing index finger at audience]… it seems rude… would be good to have other suggestions [for gestures] here” (P4). To address this, we enable users to customize the assigned gestures by selecting from two alternative gestures retrieved by the gesture identification model.
@@ 5.3.1. Final Interface
6. Implementation Details
To train the emphasis proposal model, we use Llama-2 7B [^1] https://huggingface.co/meta-llama/Llama-2-7b as the base LLM and integrate LoRA adapters with $r=128,\alpha=256$ for fine-tuning using the PEFT [^2] https://huggingface.co/docs/peft/en/index library. The base LLM remains frozen, and only the LoRA parameters are trained. Training is conducted for 20 epochs with a learning rate of $032e-$ 03 using a single NVIDIA H100 GPU.
We use the bge-base-en-v1.5-gguf [^3] https://huggingface.co/CompendiumLabs/bge-base-en-v1.5-gguf embedding model in the filtering and gesture identification module. We retrieve the top-3 gesture candidates using the established approximate K nearest neighbor search with cosine score as the similarity measure  (Kushilevitz et al., 1998). The system prompts for LLMs in both modules can be found in Appendix  B.
The gesture video clips in the database were created by enacting the gestures while wearing a Quest Pro headset, with the motion mirrored onto a humanoid avatar using the Movement SDK [^4] https://developers.meta.com/horizon/documentation/unity/move-overview/. In total, we recorded 343 clips, each 1–3 seconds long, at 1080p resolution and 30 frames per second.
The real-time speech tracking module is implemented in Python using the RealtimeSTT library [^5] https://github.com/KoljaB/RealtimeSTT for live speech-to-text conversion. It aligns spoken words with the presenter’s notes using a lexical similarity matching algorithm. Each time the rehearse button is pressed, the module resets the speech flow index to the start of the current slide’s notes. During tracking, each newly recognized word prompts a comparison between the last three spoken words and all three-word sequences within a sliding window of the speech notes. The window spans two words before to ten words after the current speech flow index.
@@ 6. Implementation Details
$\sigma=\left(1-\frac{\text{Levenshtin\ Distance}(s_{1},s_{2})}{\lvert s_{1}\rvert+\lvert s_{2}\rvert}\right)\times 100$
@@ 7.1.2. Metrics
  • Semantic matching: This is a more relaxed matching where predicted gesture regions are considered true positives if there is a 75 percent match in similarity between its vector embedding and that of a ground truth. This is an important metric because the LLM is a generative model and the generated predictions could have minor variations from the original text in terms of spelling, missing letters or punctuations. Gesture regions that are not found in the text are considered invalid and not counted as false positives, as the system directly discards them.
@@
Precision Recall F1 No. of Predicted gesture regions Length of Predictions (No. of words)
@@
GPT-4o111accessed on 21-05-2024 0.034 0.039 0.043 0.047 0.038 0.043 4.18 ±$\pm{\scriptstyle{1.67}}$ 15.41 ±$\pm{\scriptstyle{8.52}}$
Llama-2 7B 0.057 0.081 0.062 0.085 0.059 0.083 4.1 ±$\pm{\scriptstyle{1.95}}$ 8.20 ±$\pm{\scriptstyle{6.52}}$
Our Model 0.159 0.206 0.185 0.233 0.171 0.219 4.45 ±$\pm{\scriptstyle{1.26}}$ 2.92 ±$\pm{\scriptstyle{1.42}}$
Table 1. Model performance comparison for held-out test samples using direct matching (DM) and semantic matching (SM) schemes.
[…]
The table shows the precision, recall, and F1 scores for the baselines and our model for held-out samples in the database. Our model has higher scores than the baselines. It also shows average no. of predicted gesture regions in each model and the length of predictions in number of words. Our model predicts similar number of gesture regions (about 4) as off-the-shelf models but is more precise in its predictions with regions being upto 3 words long on average.
Precision Recall F1 No. of Predicted gesture regions Length of Predictions (No. of words)
@@
GPT-4o111accessed on 21-05-2024 0.038 0.038 0.002 0.002 0.028 0.028 4.48 ±$\pm{\scriptstyle{1.13}}$ 16.40 ±$\pm{\scriptstyle{9.44}}$
Llama-2 7B 0.190 0.230 0.101 0.114 0.132 0.152 4.17 ±$\pm{\scriptstyle{2.21}}$ 6.03 ±$\pm{\scriptstyle{5.02}}$
Our Model 0.348 0.490 0.171 0.233 0.230 0.316 3.86 ±$\pm{\scriptstyle{1.25}}$ 2.93 ±$\pm{\scriptstyle{1.39}}$
Table 2. Model performance comparison for unseen TED talk video using direct matching (DM) and semantic matching (SM) schemes.
[…]
The table shows the precision, recall, and F1 scores for the baselines and our model for an unseen TED video. Our model has higher scores than the baselines. It also shows average no. of predicted gesture regions in each model and the length of predictions in number of words. Our model predicts similar number of gesture regions (about 4) as off-the-shelf models but is more precise in its predictions with regions being upto 3 words long on average.
Figure 5. Distribution of users’ ratings for the appropriateness of false positives and the suitability of gestures selected by the model. Over half of false positives were rated as valid (avg. ratings ¿5), and 40% as neutral (avg. ratings 3.5–5). None were clearly inappropriate (avg. ratings ¡ 3), showing the model’s recommendations are sensible in general.
[…]
A stacked 100% bar graph showing the distribution of ratings for false positive gesture regions. Users ratings for the regions are highly positive along the dimensions of region appropriateness and the semantic gesture suitability, i.e., 60-70% of the samples have ratings above 5 on a 7 point Likert scale in both dimensions.

### v2

@@ Abstract.
Gesture, LLM, Generative AI, Presentation Talks
journalyear: 2025copyright: ccconference: The 38th Annual ACM Symposium on User InterfaceSoftware and Technology; September 28-October 1, 2025; Busan, Republic of Koreabooktitle: The 38th Annual ACM Symposium on User Interface Software and Technology (UIST ’25), September 28-October 1, 2025, Busan, Republic of Koreadoi: 10.1145/3746059.3747705isbn: 979-8-4007-2037-6/2025/09ccs: Human-centered computing Interactive systems and tools
Figure 1. GestureCoach guides speakers to perform gestures while rehearsing their talk. A gesture recommendation model predicts text segments in the presenter notes that should be emphasized with gestures and retrieves relevant semantic gestures for each segment. During rehearsal, the system highlights the segments and proactively cues a video clip of the gesture by tracking users’ speech, allowing them to integrate the gesture smoothly into their talk. Hovering over a segment allows users to preview and modify the associated gesture with alternate suggestions from the model.:
@@ 1. Introduction
As a step towards tackling these challenges, we propose GestureCoach, the first system capable of providing targeted guidance to perform gestures during rehearsal. GestureCoach combines a gesture recommendation system with a rehearsal interface that delivers real-time gesture cues, helping speakers naturally integrate gestures while speaking. The gesture recommendation system utilizes recent advances in AI chaining, by combining two specialized Large Language Model (LLM) modules, each targeting a specific aspect of gesture-based emphasis. First, the emphasis proposal module takes presenter notes as input and determines appropriate gesture regions. For this, we leverage the foundational knowledge of LLMs and further fine-tune it with annotated data from 10 TED talk videos of expert speakers. Second, the gesture identification module selects the specific semantic gesture to emphasize each predicted gesture region in a data-driven manner, retrieving relevant examples from a curated database of gestures performed by experts using a Retrieval-Augmented Generation (RAG) framework  (Lewis et al., 2020).
@@ 2. Related Work
Figure 2. Overview of the GestureCoach system. Users’ presenter notes are processed by a gesture recommendation model in the backend. A frontend rehearsal interface displays the recommended gesture regions and gestures. During rehearsal, users’ speech is tracked and gesture cues are proactively delivered to provide real-time gesture guidance. :
@@ 2.1. Co-speech Gestures for Communication
Co-speech gestures are hand and arm movements that naturally accompany speech and play a vital role in communication (Kendon, 1972; McNeill, 1992; Shattuck-Hufnagel and Prieto, 2019; Maricchiolo et al., 2020). These gestures are broadly classified into four types: beat, iconic, metaphoric, and deictic gestures (McNeill, 1992). Beat gestures are rhythmic movements that align with speech prosody to emphasize its structure. In contrast, iconic, metaphoric, and deictic gestures are semantic gestures that are directly related to the linguistic content being spoken, enhancing content delivery during communication  (Maricchiolo et al., 2020; Kendon, 1994; Hostetter, 2011). Iconic gestures physically resemble the objects or actions they represent, such as mimicking an object’s shape or reenacting a movement. Metaphoric gestures convey abstract concepts through spatial metaphors, like cupped hands symbolizing ”holding” an idea. Deictic gestures involve pointing to indicate referents, whether physically, such as gesturing toward an object, or abstractly, such as indicating an imaginative space.
Semantic gestures, which are the focus of this study, have been empirically shown to significantly benefit listener comprehension, especially when they depict motor actions or convey non-redundant information (Hostetter, 2011). These gestures also help speakers express their intentions through complementary movements, making the message clearer and more engaging for listeners (McNeill, 1992; Kendon, 2004; Hinnell, 2019). For example, moving one’s hands from side to side can indicate contrast when saying “on the other hand” (Hinnell, 2019). Moreover, encouraging speakers to use hand gestures has been found to increase narrative length and influence prosodic features, such as fundamental frequency and intensity, thereby improving intonation (Cravotta et al., 2019). These findings show that semantic gestures play an important role in making communication and presentations more effective.
@@ 2.3. Modeling Semantic Gestures
In this work, we aim to disentangle the tasks of predicting when to gesture and what semantic gesture to perform (Neff, 2016). Linguistic research suggests that semantic gestures often co-occur with specific spoken discourse markers, such as discourse connectives (e.g., “however,” “for example”) that indicate topic shifts, and stance markers (e.g., “absolutely” “probably”) that reflect the speaker’s level of certainty (Mcneill et al., 2014; Laparle, 2021; Calbris, 2011; Suresh et al., 2025). For instance, raised index fingers tend to co-occur with contrastive discourse markers like ”but,” ”however,” that signal exceptions (Inbar, 2022). These insights suggest that there are detectable patterns that can help determine when to gesture directly from text, which we address by fine-tuning an LLM to predict these gesture regions. Determining what semantic gesture to perform, on the other hand, is more challenging as there are no clear rules on what gesture best fits a given context. Hence, we guide our modeling using a data-driven approach, dynamically retrieving contextually appropriate gestures from a corpus of expert gestures using RAG.
@@ 4. GestureCoach: Recommendation Model
Figure 3. Gesture recommendation model architecture. It consists of two LLM-based modules: 1) The emphasis proposal module is fine-tuned on data from expert speakers to predict gesture regions in the presenter notes. The predicted regions are filtered and passed to 2) the gesture identification module which selects the most suitable semantic gesture from the gesture database using Retrieval-Augmented Generation.:
@@ 4.1. Gesture Database
To create a gesture database, we first collected TED talk videos of expert speakers who performed semantic gestures. The database consists of annotated gesture regions in the talk transcripts where the speaker performed semantic gestures. Each gesture region is also associated with a video clip of a humanoid avatar performing the semantic gesture.
@@
The annotated gesture regions have an average length of 3.10 (± 2.48) words with an average of 3.81 (± 2.38) gestures regions in a sample. The gesture video clips in the database were created by enacting the gestures while wearing a Quest Pro headset, with the motion mirrored onto a humanoid avatar using the Movement SDK [^1] https://developers.meta.com/horizon/documentation/unity/move-overview/. In total, we recorded 343 clips, each 1–3 seconds long, at 1080p resolution and 30 frames per second.
@@ 4.2.1. Fine-tuning with Data Augmentation
We use Llama-2 7B [^2] https://huggingface.co/meta-llama/Llama-2-7b as the base LLM and integrate LoRA adapters with $r=128,\alpha=256$ for fine-tuning using the PEFT [^3] https://huggingface.co/docs/peft/en/index library. The base LLM remains frozen, and only the LoRA parameters are trained. Training is conducted for 20 epochs with a learning rate of 2e032e-03 using a single NVIDIA H100 GPU. For fine tuning and subsequent inference, we pass the input text using the standard prompt format [^4] https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-2/ (seen in Figure  3) of Llama 2 models.
@@
We include a filtering module that ensures predictions from the emphasis proposal module align with the original presenter notes before passing them to the gesture identification module. This step is important because the generative nature of LLMs can lead to predictions where the gesture regions do not match the original text verbatim. Such differences can range from minor ones (e.g., a missing apostrophe) to complete paraphrasing. The filtering process first attempts a verbatim match between the predicted gesture region and the presenter notes. If no exact match is found, we apply a sliding window approach to search for the most similar gesture region in the notes. Specifically, we compute the cosine similarity between the embedding of the predicted gesture region and the embeddings of each text window in the presenter notes. We set a similarity threshold of 75%, above which a match is accepted. If no sufficiently similar match is found, the prediction is discarded. We used bge-base-en-v1.5-gguf [^5] https://huggingface.co/CompendiumLabs/bge-base-en-v1.5-gguf as the embedding model.
@@ 4.3. Gesture Identification
We developed a custom RAG framework using the bge-base-en-v1.5-gguf model to get semantic vector embeddings, FAISS [^6] https://github.com/facebookresearch/faiss to create and efficiently search over the vector database, and the GPT-4o model to generate the final gesture candidate. We retrieve the top-3 gesture candidates using the established approximate K nearest neighbor search with cosine score as the similarity measure  (Kushilevitz et al., 1998). The prompts for LLMs in both modules can be found in Appendix  B.
@@ 5. GestureCoach: Rehearsal Interface
Figure 4. Early Wizard-of-Oz prototype that simulates a real-time gesture cueing interface for rehearsing talks.:
@@ 5.2. Pilot Study
We recruited 5 participants (3 males, 2 females, aged 25 to 45 years) who volunteered for the study. Four of them self-identified as novice speakers who had given 1-2 formal talks in the past and one as an expert who had given more than 10 talks. After an introduction and walkthrough of the prototype, participants were asked to rehearse for a talk, emphasizing the highlighted regions in the presenter notes by mimicking gesture cues displayed in the gesture window.
@@ 5.3. Design Insights & Final Interface
  • Gesture cue visualization. Early in our design process, we chose to demonstrate the gestures rather than representing them symbolically. While lightweight symbols such as emojis or gesture icons may be quicker for users to interpret during rehearsal, it can also be challenging to convey the spatial relationships and movement dynamics of gestures with such static representations. Our pilots indicate that users were generally comfortable following gesture demonstrations presented in a third-person, upper-body view. This perspective allowed them to clearly observe arm and hand movements while understanding their spatial relation to the upper body, aiding effective gesturing. However, two participants reported that guidance from a real human felt distracting, inducing a sense of social evaluation, particularly due to eye contact and facial expressions: ”Its a bit embarrassing …I feel like someone’s looking at me… it make me more self-aware and affects my practice” (P2). Hence, in our final interface we use a animated humanoid avatar (upper-body only) captured from below the face, focusing on arm and hand motions to emphasize the gesture.
@@
  • Modifying predicted gestures. Users occasionally found that some of the gesture regions or the suggested gesture did not align with their expectations due to individual communication styles or cultural differences in gesturing norms. Certain gestures may carry different meanings across cultures, making them less suitable or even inappropriate for some users: ”I don’t point like this [pointing index finger at audience]… it seems rude… would be good to have other suggestions [for gestures] here” (P4). To address this, we enable users to customize the assigned gestures by selecting from two alternative gestures retrieved by the gesture identification model.
@@ 5.3.1. Final Interface
The real-time speech tracking module is implemented in Python using the RealtimeSTT library [^7] https://github.com/KoljaB/RealtimeSTT for live speech-to-text conversion. It aligns spoken words with the presenter’s notes using a lexical similarity matching algorithm. Each time the rehearse button is pressed, the module resets the speech flow index to the start of the current slide’s notes. During tracking, each newly recognized word prompts a comparison between the last three spoken words and all three-word sequences within a sliding window of the speech notes. The window spans two words before and ten words after the current speech flow index.
@@
$\sigma=\left(1-\frac{\text{Levenshtein\ Distance}(s_{1},s_{2})}{\lvert s_{1}\rvert+\lvert s_{2}\rvert}\right)\times 100$
@@ 6.1.2. Metrics
  • Semantic matching: This is a more relaxed matching where predicted gesture regions are considered true positives if there is a 75 percent match in similarity between its vector embedding and that of a ground truth. This is an important metric because the LLM is a generative model and the generated predictions could have minor variations from the original text in terms of spelling, missing letters or punctuation. Gesture regions that are not found in the text are considered invalid and not counted as false positives, as the system directly discards them.
@@
Precision Recall F1
No. of Predicted
gesture regions
Length of Predictions
(No. of words)
@@
GPT-4o111accessed on 21-05-2024 0.034 0.039 0.043 0.047 0.038 0.043 4.18 $\pm{\scriptstyle{1.67}}$ 15.41 $\pm{\scriptstyle{8.52}}$
Llama-2 7B 0.057 0.081 0.062 0.085 0.059 0.083 4.1 $\pm{\scriptstyle{1.95}}$ 8.20 $\pm{\scriptstyle{6.52}}$
Our Model 0.159 0.206 0.185 0.233 0.171 0.219 4.45 $\pm{\scriptstyle{1.26}}$ 2.92 $\pm{\scriptstyle{1.42}}$
Table 1. Model performance comparison for held-out test samples using direct matching (DM) and semantic matching (SM) schemes.:
Precision Recall F1
No. of Predicted
gesture regions
Length of Predictions
(No. of words)
@@
GPT-4o111accessed on 21-05-2024 0.038 0.038 0.002 0.002 0.028 0.028 4.48 $\pm{\scriptstyle{1.13}}$ 16.40 $\pm{\scriptstyle{9.44}}$
Llama-2 7B 0.190 0.230 0.101 0.114 0.132 0.152 4.17 $\pm{\scriptstyle{2.21}}$ 6.03 $\pm{\scriptstyle{5.02}}$
Our Model 0.348 0.490 0.171 0.233 0.230 0.316 3.86 $\pm{\scriptstyle{1.25}}$ 2.93 $\pm{\scriptstyle{1.39}}$
Table 2. Model performance comparison for unseen TED talk video using direct matching (DM) and semantic matching (SM) schemes.:
Figure 5. Distribution of users’ ratings for the appropriateness of false positives and the suitability of gestures selected by the model. Over half of false positives were rated as valid (avg. ratings ¿5), and 40% as neutral (avg. ratings 3.5–5). None were clearly inappropriate (avg. ratings ¡ 3), showing the model’s recommendations are sensible in general.: